Section 6

Topology

In case of several authors, Invited Speakers are marked with a ^{∗}.

A. N. Dranishnikov: Dimension Theory and Large Riemannian Manifolds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 423

W. G. Dwyer: Lie Groups and p-Compact Groups . . . . . . . . . . . . . . . . . II 433

Ronald Fintushel^{∗} and Ronald J. Stern^{∗}: Constructions of Smooth 4-Manifolds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 443

Michael H. Freedman: Topological Views on Computational Complexity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 453

Mark Mahowald: Toward a Global Understanding of π_{∗}(S^{n}) . . . . . . II 465

Tomotada Ohtsuki: A Filtration of the Set of Integral Homology 3-Spheres . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 473

Bob Oliver: Vector Bundles over Classifying Spaces . . . . . . . . . . . . . . . . II 483

Clifford Henry Taubes: The Geometry of the Seiberg-Witten Invariants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 493

Documenta Mathematica · Extra Volume ICM 1998 · II ·

422

Documenta Mathematica · Extra Volume ICM 1998 · II ·